# Hengwei Bian1,2,    Lingdong Kong1,3   Haozhe Xie4   Liang Pan1,†,‡   Yu Qiao1   Ziwei Liu4

### v2

@@ 4.2 Diffusion Transformer for HexPlane
After training the VAE, 4D semantic scenes can be embedded as HexPlane $\mathcal{H}=[\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},\mathcal{P}_{tz}]$. Building upon $\mathcal{H}$, we aim to leverage a DiT (Peebles & Xie, 2023) model $D_{\tau}$ to generate novel HexPlane, which could be further decoded as novel 4D scenes (see Fig. 2(b)). However, training a DiT using token sequences naively generated from each feature plane of HexPlane could not guarantee high generation quality, mainly due to the absence of modeling spatial and temporal relations among the tokens.
Padded Rollout Operation. Given that the feature planes of HexPlane may share spatial or temporal dimensions, we employ the Padded Rollout Operation (PRO) to systematically arrange all six planes into a unified square feature map, incorporating zero paddings in the uncovered corner areas. As shown in Fig. 5, the dimension of the 2D square feature map is $(\frac{X}{d_{X}}+\frac{Z}{d_{Z}}+\frac{T}{d_{T}})$, which minimizes the area for padding, where $d_{X}$, $d_{Z}$, and $d_{T}$ represent the downsampling rates along the X, Z, and T axes, respectively. Subsequently, we follow DiT to first “patchify” the constructed 2D feature map, converting it into a sequence of $N=((\frac{X}{d_{X}}+\frac{Z}{d_{Z}}+\frac{T}{d_{T}})/p)^{2}$ tokens, where $p$ is the patch size, chosen so each token holds information from one feature plane. Following patchification, we apply the frequency-based positional embeddings to all tokens similar to DiT. Note that tokens corresponding to padding areas are excluded from the diffusion process. Consequently, the proposed PRO offers an efficient method for modeling spatial and temporal relationships within the token sequence.
Conditional Generation. DiT enables conditional generation through the use of Classifier-Free Guidance (CFG) (Ho & Salimans, 2022). To incorporate conditions into the generation process, we designed two branches for condition insertion (see Fig. 5). For any condition $c$, we use the adaLN-Zero technique from DiT, generating scale and shift parameters from $c$ and injecting them before and after the attention and feed-forward layers. To handle the complexity of image-based conditions, we add a cross-attention block to better integrate the image condition into the DiT block.
@@ 4.3 Downstream Applications
Inspired by SemCity (Lee et al., 2024), we also manipulate the HexPlane during sampling to: 5) Inpaint: Edit 4D scenes by masking HexPlane regions and guiding sampling with the masked areas. For more applications and implementation details, kindly refer to Sec. A.5 in the Appendix.
Table 1: Comparisons of 4D Scene Reconstruction. We report the mIoU scores of OccSora (Wang et al., 2024) and our DynamicCity framework on the CarlaSC, Occ3D-Waymo, and Occ3D-nuScenes datasets, respectively, under different resolutions and sequence lengths. Symbol denotes score reported in the OccSora paper. Other scores are reproduced using the official code.
@@
(Wang et al., 2024) (DynamicCity)
CarlaSC (Wilson et al., 2022) 10 128$\times$128$\times$8 4 41.01% 79.61% (+38.6%)
10 128$\times$128$\times$8 8 39.91% 76.18% (+36.3%)
10 128$\times$128$\times$8 16 33.40% 74.22% (+40.8%)
10 128$\times$128$\times$8 32 28.91% 59.31% (+30.4%)
Occ3D-Waymo (Tian et al., 2023) 9 200$\times$200$\times$16 16 36.38% 68.18% (+31.8%)
Occ3D-nuScenes (Tian et al., 2023) 11 200$\times$200$\times$16 16 13.70% 56.93% (+43.2%)
11 200$\times$200$\times$16 32 13.51% 42.60% (+29.1%)
17 200$\times$200$\times$16 32 13.41% 40.79% (+27.3%)
17 200$\times$200$\times$16 32 27.40% 40.79% (+13.4%)
Table 2: Comparisons of 4D Scene Generation. We report the Inception Score (IS), Fréchet Inception Distance (FID), Kernel Inception Distance (KID), and the Precision (P) and Recall (R) rates of OccSora (Wang et al., 2024) and our DynamicCity framework on the CarlaSC and Occ3D-Waymo datasets, respectively, in both the 2D and 3D spaces.
@@
5 Experiments
5.1 Experimental Details
Datasets. We train the proposed model on the 1Occ3D-Waymo, 2Occ3D-nuScenes, and 3CarlaSC datasets. The former two from Occ3D (Tian et al., 2023) are derived from Waymo (Sun et al., 2020) and nuScenes (Caesar et al., 2020), where LiDAR point clouds have been completed and voxelized to form occupancy data. Each occupancy scene has a resolution of $200\times 200\times 16$, covering a region centered on the ego vehicle, extending $40$ meters in all directions and $6.4$ meters vertically. The CarlaSC dataset (Wilson et al., 2022) is a synthetic occupancy dataset, with a scene resolution of $128\times 128\times 8$, covering a region $25.6$ meters around the ego vehicle, with a height of $3$ meters.
Implementation Details. Our experiments are conducted using eight NVIDIA A100-80G GPUs. The global batch size used for training the VAE is $8$, while the global batch size for training the DiT is $128$. Our latent HexPlane $\mathcal{H}$ is compressed to half the size of the input $\mathbf{Q}$ in each dimension, with the latent channels $C=16$. The weight for the Lovász-softmax and KL terms are set to $1$ and $0.005$, respectively. The learning rate for the VAE is $10^{-3}$, while the learning rate for the DiT is $10^{-4}$.
Evaluation Metrics. The mean intersection over union (mIoU) metric is used to evaluate the reconstruction results of VAE. For DiT, Inception Score, FID, KID, Precision, and Recall are calculated for evaluation. Specifically, we follow prior work (Lee et al., 2024; Wang et al., 2024) by rendering 3D scenes into 2D images and utilizing conventional 2D evaluation pipelines for assessment. Additionally, we train the 3D Encoder to directly extract features from the 3D data and calculate the metrics. For more details, kindly refer to Sec. A.2 in the Appendix.
Figure 6: Dynamic Scene Generation Results. We provide unconditional generation scenes from the $1$st, $8$th, and $16$th frames on Occ3D-Waymo (Left) and CarlaSC (Right), respectively. Kindly refer to the Appendix for complete sequential scenes and longer temporal modeling examples.
5.2 4D Scene Reconstruction & Generation
Reconstruction. To evaluate the effectiveness of the proposed VAE in encoding the 4D occupancy sequence, we compare it with OccSora (Wang et al., 2024) using the CarlaSC, Occ3D-Waymo, and Occ3D-nuScenes datasets. As shown in Tab. 1, DynamicCity outperforms OccSora on these datasets, achieving mIoU improvements of 38.6%, 31.8%, and 43.2% respectively, when the input number of frames is 16. These results highlight the superior performance of the proposed VAE.
Generation. To demonstrate the effectiveness of DynamicCity in 4D scene generation, we compare the generation results with OccSora (Wang et al., 2024) on the Occ3D-Waymo and CarlaSC datasets. As shown in Tab. 2, the proposed method outperforms OccSora in terms of perceptual metrics in both 2D and 3D spaces. These results show that our model excels in both generation quality and diversity. Fig. 6 and Fig. 16 show the 4D scene generation results, demonstrating that our model is capable of generating large dynamic scenes in both real-world and synthetic datasets. Our model not only exhibits the ability to generate moving scenes with static semantics shifting as a whole, but it is also capable of generating dynamic elements such as vehicles and pedestrians.
Figure 7: Dynamic Scene Generation Applications. We demonstrate the capability of our model on a diverse set of downstream tasks. We show the $1$st, $8$th, and $16$th frames for simplicity. Kindly refer to the Appendix for complete sequential scenes and longer temporal modeling examples.
5.3 Ablation Studies
Table 3: Ablation Study on VAE Network Structures. We report the mIoU scores, training time (second-per-iteration), and training-time memory consumption (VRAM) of different Encoder and Decoder configurations on CarlaSC and Occ3D-Waymo, respectively. Note that “ESS” denotes “Expansion & Squeeze”. The best and second-best values are in bold and underlined.
@@ D.S. Rates CarlaSC Occ3D-Waymo
[…]
@@ 6 Conclusion
Acknowledgments
This study is supported by the National Key R&\&&D Program of China No.2022ZD0160102. This study is also supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012, MOE-T2EP20223-0002), NTU NAP, and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s). This study is also under the programme DesCartes and is supported by the National Research Foundation, Prime Minister’s Office, Singapore, under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.
We sincerely thank our reviewers and ACs for the time and effort they devoted during the review.
@@ Acknowledgments
[…]
Appendix A A Additional Implementation Details
@@
A.1 Datasets
Our experiments primarily utilize two datasets: Occ3D-Waymo (Tian et al., 2023) and CarlaSC (Wilson et al., 2022). Additionally, we also evaluate our VAE on Occ3D-nuScenes (Tian et al., 2023).
The Occ3D-Waymo dataset is derived from real-world Waymo Open Dataset (Sun et al., 2020) data, where occupancy sequences are obtained through multi-frame fusion and voxelization processes. Similarly, Occ3D-nuScenes is generated from the real-world nuScenes (Caesar et al., 2020) dataset using the same fusion and voxelization operations. On the other hand, the CarlaSC dataset is generated from simulated scenes and sensor data, yielding occupancy sequences.
@@ A.1 Datasets
Table 6: Summary of Semantic Label Mappings. We unify the semantic classes between CarlaSC (Wilson et al., 2022), Occ3D-Waymo (Tian et al., 2023), and Occ3D-nuScenes (Tian et al., 2023) datasets for semantic scene generation.
@@ A.2 DiT Evaluation Metrics
$\text{IS}=\exp\left(\mathbb{E}_{\mathbf{Q}\sim p_{g}}\left[D_{\mathrm{KL}}(p(y|\mathbf{Q})\parallel p(y))\right]\right)~{},$ (5)
[…]
$D_{\mathrm{KL}}(p(y|\mathbf{Q})\parallel p(y))=\sum_{i}p(y_{i}|\mathbf{Q})\log\frac{p(y_{i}|\mathbf{Q})}{p(y_{i})}~{}.$ (6)
@@
$\text{FID}=\|\mu_{r}-\mu_{g}\|^{2}+\mathrm{Tr}\left(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2}\right)~{},$ (7)
@@
$\text{KID}=\text{MMD}^{2}(\phi(\mathbf{Q}_{r}),\phi(\mathbf{Q}_{g}))~{},$ (8)
@@
$\text{MMD}^{2}(X,Y)=\frac{1}{m(m-1)}\sum_{i\neq j}k(x_{i},x_{j})+\frac{1}{n(n-1)}\sum_{i\neq j}k(y_{i},y_{j})-\frac{2}{mn}\sum_{i,j}k(x_{i},y_{j})~{},$ (9)
@@
$\text{Precision}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left((\mathbf{f}_{g}-\mu_{r})^{\top}\Sigma_{r}^{-1}(\mathbf{f}_{g}-\mu_{r})\leq\chi^{2}\right)~{},$ (10)
@@
$\text{Recall}=\frac{1}{M}\sum_{j=1}^{M}\mathbb{I}\left((\mathbf{f}_{r}-\mu_{g})^{\top}\Sigma_{g}^{-1}(\mathbf{f}_{r}-\mu_{g})\leq\chi^{2}\right)~{},$ (11)
@@
2D Evaluations. We render 3D scenes as 2D images for 2D evaluations. To ensure fair comparisons, we use the same semantic colormap and camera settings across all experiments. Fig. 8 shows an example of a rendered 2D semantic colormap. We use an InceptionV3 (Szegedy et al., 2015) model to compute the Inception Score (IS), Fréchet Inception Distance (FID), and Kernel Inception Distance (KID) scores, while Precision and Recall are computed using a VGG-16 (Simonyan & Zisserman, 2015) model. We train both 2D backbones using semantic colormap data.
@@
3D Evaluations. For 3D data, we train a MinkowskiUNet (Choy et al., 2019) as an autoencoder. We adopt the latest implementation from SPVNAS (Tang et al., 2020), which supports optimized sparse convolution operations. The features were extracted by applying average pooling to the output of the final downsampling block.
A.3 Model Details
General Training Details. We implement both the VAE and DiT models using PyTorch (Paszke et al., 2019). We utilize PyTorch’s mixed precision and replace all attention mechanisms with FlashAttention (Dao et al., 2022) to accelerate training and reduce memory usage. AdamW is used as the optimizer for all models.

### v3

@@ 4.2 Diffusion Transformer for HexPlane
After training the VAE, 4D semantic scenes can be embedded as HexPlane $\mathcal{H}=[\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},\mathcal{P}_{tz}]$. Building upon $\mathcal{H}$, we aim to leverage a DiT [26] model $D_{\tau}$ to generate novel HexPlane, which could be further decoded as novel 4D scenes (see Fig. 2(b)).
However, training a DiT using token sequences naively generated from each feature plane of HexPlane could not guarantee high generation quality, mainly due to the absence of modeling spatial and temporal relations among the tokens.
Padded Rollout Operation. Given that the feature planes of HexPlane may share spatial or temporal dimensions, we employ the Padded Rollout Operation (PRO) to systematically arrange all six planes into a unified square feature map, incorporating zero paddings in the uncovered corner areas. As shown in Fig. 5, the dimension of the 2D square feature map is $(\frac{X}{d_{X}}+\frac{Z}{d_{Z}}+\frac{T}{d_{T}})$, which minimizes the area for padding, where $d_{X}$, $d_{Z}$, and $d_{T}$ represent the downsampling rates along the X, Z, and T axes, respectively.
Subsequently, we follow DiT to first “patchify” the constructed 2D feature map, converting it into a sequence of $N=((\frac{X}{d_{X}}+\frac{Z}{d_{Z}}+\frac{T}{d_{T}})/p)^{2}$ tokens, where $p$ is the patch size, chosen so each token holds information from one feature plane. Following patchification, we apply the frequency-based positional embeddings to all tokens similar to DiT. Note that tokens corresponding to padding areas are excluded from the diffusion process. Consequently, the proposed PRO offers an efficient method for modeling spatial and temporal relationships within the token sequence.
Conditional Generation. DiT enables conditional generation through the use of Classifier-Free Guidance (CFG) [12]. To incorporate conditions into the generation process, we designed two branches for condition insertion (see Fig. 5). For any condition $c$, we use the adaLN-Zero technique from DiT, generating scale and shift parameters from $c$ and injecting them before and after the attention and feed-forward layers. To handle the complexity of image-based conditions, we add a cross-attention block to better integrate the image condition into the DiT block.
@@ 4.3 Downstream Applications
Inspired by SemCity [17], we also manipulate the HexPlane during sampling to: 5) Inpaint: Edit 4D scenes by masking HexPlane regions and guiding sampling with the masked areas. For more applications and implementation details, kindly refer to Sec. 7.5 in the Appendix.
Figure 6: Dynamic Scene Generation Results. We provide unconditional generation scenes from the 11st, 88th, and $16$th frames on Occ3D-Waymo (Left) and CarlaSC (Right), respectively. Kindly refer to the Appendix for complete sequential scenes and longer temporal modeling examples.
Table 1: Comparisons of 4D Scene Reconstruction. We report the mIoU scores of OccSora [41] and our DynamicCity framework on the CarlaSC, Occ3D-Waymo, and Occ3D-nuScenes datasets, respectively, under different resolutions and sequence lengths. Symbol denotes score reported in OccSora [41]. Other scores are reproduced using the official code.
@@
[41] (DynamicCity)
CarlaSC [42] 10 $\times128$$\times$8 4 41.01% 79.61% (+38.6%)
10 $\times128$$\times$8 8 39.91% 76.18% (+36.3%)
10 $\times128$$\times$8 16 33.40% 74.22% (+40.8%)
10 $\times128$$\times$8 32 28.91% 59.31% (+30.4%)
Occ3D-Waymo [40] 9 $\times200$$\times$16 16 36.38% 68.18% (+31.8%)
Occ3D-nuScenes [40] 11 $\times200$$\times$16 16 13.70% 56.93% (+43.2%)
11 $\times200$$\times$16 32 13.51% 42.60% (+29.1%)
17 $\times200$$\times$16 32 13.41% 40.79% (+27.3%)
17 $\times200$$\times$16 32 27.40% 40.79% (+13.4%)
Table 2: Comparisons of 4D Scene Generation. We report the Inception Score (IS), Fréchet Inception Distance (FID), Kernel Inception Distance (KID), and the Precision (P) and Recall (R) rates of OccSora [41] and our DynamicCity framework on the CarlaSC and Occ3D-Waymo datasets, respectively, in both the 2D and 3D spaces.
Dataset Method #Frames $^{{\color[rgb]{0.890625,0.421875,0.76171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.890625,0.421875,0.76171875}\textbf{2D}}} Metric$$^{{\color[rgb]{0.5,0.625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.625,0.8828125}\textbf{3D}}}$
@@
Table 3: Ablation Study on VAE Network Structures. We report mIoU, training time (seconds-per-iteration), and training-time memory consumption (VRAM) of different Encoder and Decoder configurations on CarlaSC and Occ3D-Waymo. Note that “ESS” denotes “Expansion & Squeeze”. The best and second-best values are in bold and underlined.
@@
5 Experiments
5.1 Experimental Details
Datasets. We train the proposed model on the 1Occ3D-Waymo, 2Occ3D-nuScenes, and 3CarlaSC datasets. The former two from Occ3D [40] are derived from Waymo [37] and nuScenes [6], where LiDAR point clouds have been completed and voxelized to form occupancy data. Each occupancy scene has a resolution of $200\times 200\times 16$, covering a region centered on the ego vehicle, extending $40$ meters in all directions and $6.4$ meters vertically. The CarlaSC dataset [42] is a synthetic occupancy dataset, with a scene resolution of $128\times 128\times 8$, covering a region $25.6$ meters around the ego vehicle, with a height of $3$ meters.
Implementation Details. Our experiments are conducted using eight NVIDIA A100-80G GPUs. The global batch size used for training the VAE is $8$, while the global batch size for training the DiT is $128$. Our latent HexPlane $\mathcal{H}$ is compressed to half the size of the input $\mathbf{Q}$ in each dimension, with the latent channels $C=16$. The weights for the Lovász-softmax and KL terms are set to $1$ and $0.005$, respectively. The learning rate for the VAE is $10^{-3}$, while the learning rate for the DiT is $10^{-4}$.
Figure 7: Dynamic Scene Generation Applications. We demonstrate the capability of our model on a diverse set of downstream tasks. We show the 11st, 88th, and $16$th frames for simplicity. Kindly refer to the Appendix for complete sequential scenes and longer temporal modeling examples.
Evaluation Metrics. The mean intersection over union (mIoU) metric is used to evaluate the reconstruction results of VAE. For DiT, Inception Score, FID, KID, Precision, and Recall are calculated for evaluation. Specifically, we follow prior work [17, 41] by rendering 3D scenes into 2D images and utilizing conventional 2D evaluation pipelines for assessment. Additionally, we train the 3D Encoder to directly extract features from the 3D data and calculate the metrics. For more details, kindly refer to Sec. 7.2 in the Appendix.
5.2 4D Scene Reconstruction & Generation
Reconstruction. To evaluate the effectiveness of the proposed VAE in encoding the 4D occupancy sequence, we compare it with OccSora [41] using the CarlaSC, Occ3D-Waymo, and Occ3D-nuScenes datasets. As shown in Tab. 1, DynamicCity outperforms OccSora on these datasets, achieving mIoU improvements of 38.6%, 31.8%, and 43.2%, respectively, when the input number of frames is 16. These results highlight the superior performance of the proposed VAE.
Generation. To demonstrate the effectiveness of DynamicCity in 4D scene generation, we compare the generation results with OccSora [41] on the Occ3D-Waymo and CarlaSC datasets. As shown in Tab. 2, the proposed method outperforms OccSora in terms of perceptual metrics in both 2D and 3D spaces. These results show that our model excels in both generation quality and diversity. Fig. 6 and Fig. 16 show the 4D scene generation results, demonstrating that our model is capable of generating large dynamic scenes in both real-world and synthetic datasets. Our model not only exhibits the ability to generate moving scenes with static semantics shifting as a whole, but it is also capable of generating dynamic elements such as vehicles and pedestrians.
5.3 Ablation Study
@@ D.S. Rates CarlaSC Occ3D-Waymo
Method $^{{\color[rgb]{0.890625,0.421875,0.76171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.890625,0.421875,0.76171875}\textbf{2D}}} Metric$$^{{\color[rgb]{0.5,0.625,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.5,0.625,0.8828125}\textbf{3D}}}$
@@ 6 Conclusion
7 Additional Implementation Details
@@
7.1 Datasets
Our experiments primarily utilize two datasets: Occ3D-Waymo [40] and CarlaSC [42]. Additionally, we also evaluate our VAE on Occ3D-nuScenes [40].
The Occ3D-Waymo dataset is derived from real-world Waymo Open Dataset [37] data, where occupancy sequences are obtained through multi-frame fusion and voxelization processes. Similarly, Occ3D-nuScenes is generated from the real-world nuScenes [6] dataset using the same fusion and voxelization operations. On the other hand, the CarlaSC dataset is generated from simulated scenes and sensor data, yielding occupancy sequences.
@@ 7.1 Datasets
Table 6: Summary of Semantic Label Mappings. We unify the semantic classes between CarlaSC [42], Occ3D-Waymo [40], and Occ3D-nuScenes [40] datasets for semantic scene generation.
@@ 7.2 DiT Evaluation Metrics
$\text{IS}=\exp\left(\mathbb{E}_{\mathbf{Q}\sim p_{g}}\left[D_{\mathrm{KL}}(p(y|\mathbf{Q})\parallel p(y))\right]\right)\penalty 10000\ ,$ (5)
where $p_{g}$ represents the distribution of generated samples. $p(y|\mathbf{Q})$ is the conditional label distribution given by the Inception model for a generated sample $\mathbf{Q}$. $p(y)=\int p(y|\mathbf{Q})p_{g}(\mathbf{Q})\,d\mathbf{Q}$ is the marginal distribution over all generated samples. $D_{\mathrm{KL}}(p(y|\mathbf{Q})\parallel p(y))$ is the Kullback-Leibler divergence, defined as follows:
$D_{\mathrm{KL}}(p(y|\mathbf{Q})\parallel p(y))=\sum_{i}p(y_{i}|\mathbf{Q})\log\frac{p(y_{i}|\mathbf{Q})}{p(y_{i})}\penalty 10000\ .$ (6)
@@
$\text{FID}=\|\mu_{r}-\mu_{g}\|^{2}+\mathrm{Tr}\left(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2}\right)\penalty 10000\ ,$ (7)
@@
$\text{KID}=\text{MMD}^{2}(\phi(\mathbf{Q}_{r}),\phi(\mathbf{Q}_{g}))\penalty 10000\ ,$ (8)
@@
$\text{MMD}^{2}(X,Y)=\frac{1}{m(m-1)}\sum_{i\neq j}k(x_{i},x_{j})+\frac{1}{n(n-1)}\sum_{i\neq j}k(y_{i},y_{j})-\frac{2}{mn}\sum_{i,j}k(x_{i},y_{j})\penalty 10000\ ,$ (9)
@@
$\text{Precision}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left((\mathbf{f}_{g}-\mu_{r})^{\top}\Sigma_{r}^{-1}(\mathbf{f}_{g}-\mu_{r})\leq\chi^{2}\right)\penalty 10000\ ,$ (10)
@@
$\text{Recall}=\frac{1}{M}\sum_{j=1}^{M}\mathbb{I}\left((\mathbf{f}_{r}-\mu_{g})^{\top}\Sigma_{g}^{-1}(\mathbf{f}_{r}-\mu_{g})\leq\chi^{2}\right)\penalty 10000\ ,$ (11)
@@
2D Evaluations. We render 3D scenes as 2D images for 2D evaluations. To ensure fair comparisons, we use the same semantic colormap and camera settings across all experiments. Fig. 8 shows an example of a rendered 2D semantic colormap. We use an InceptionV3 [38] model to compute the Inception Score (IS), Fréchet Inception Distance (FID), and Kernel Inception Distance (KID) scores, while Precision and Recall are computed using a VGG-16 [34] model. We train both 2D backbones using semantic colormap data.
@@
3D Evaluations. For 3D data, we train a MinkowskiUNet [9] as an autoencoder. We adopt the latest implementation from SPVNAS [39], which supports optimized sparse convolution operations. The features were extracted by applying average pooling to the output of the final downsampling block.
7.3 Model Details
General Training Details. We implement both the VAE and DiT models using PyTorch [25]. We utilize PyTorch’s mixed precision and replace all attention mechanisms with FlashAttention [10] to accelerate training and reduce memory usage. AdamW is used as the optimizer for all models.